A man has $10,000 to invest, some in a relatively safe account
earning 5% interest per year, and the rest in more speculative
investments earning 12% per year. If the total interest earned for the
year was $955, how much was invested at each rate?

Answer by mananth(16946)

Part I 5.00% per annum ------------- Amount invested =x
Part II 12.00% per annum ------------ Amount invested = y
10000
Interest----- 955.00

Part I 5.00% per annum ---x
Part II 12.00% per annum ---y
Total investment
x + 1 y= 10000 -------------1
Interest on both investments
5.00% x + 12.00% y= 955
Multiply by 100
5 x + 12 y= 95500.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -50000.00
Add this to (2)
0 x 7 y= 45500
divide by 7
y = 6500
Part I 5.00% $ 3500
Part II 12.00% $ 6500

CHECK
3500 --------- 5.00% ------- 175.00
6500 ------------- 12.00% ------- 780.00
Total -------------------- 955.00

m.ananth@hotmail.ca